# Iterative LLM-Based Generation and Refinement of Distracting Conditions in Math Word Problems

### v1

@@ III-C Rejection Mechanism
Regeneration Prompts The question you generate does not comply with all the requirements. Considering the following explanation on your errors, revise the question again. Take the error and initial requirements into consideration. Directly give the full revised question without explanations. Q1: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … How much will each of them pay if they will split the bill equally? Q2: Five friends eat at a fast-food chain and order the following: 5 pieces of hamburger that cost $\mathdollar 3$ each … There will be additional $\mathdollar 5$ service fee when the total bill is more than $\mathdollar 26$. How much will each of them pay if they will split the bill equally? Explanation: The solution does not fit the question because the equation used in the solution is incorrect … should be 5, not 6.
@@
IV Experiment and Result
IV-A Settings
@@ IV-B Generated Dataset
As illustrated, the problems in the IGC-MWP dataset exhibit clear advantages and avoid the undesirable characteristics discussed earlier. For instance, in Problem 1 of Fig. 9, the original story describes slices of pizza, and the added condition naturally follows this theme rather than introducing unrelated context. The original question involves two pizza sizes, while the added condition introduces a medium pizza with a comparable number of slices (12). This new condition adds meaningful distraction—students must now reason about the number of slices in each pizza type and correctly distinguish between large, small, and medium sizes.
Similarly, in Problems 2 and 3 of Fig. 9, the added conditions involve the number of lines on pages or the number of pages in books, both of which are conceptually tied to the core entities of the original stories about reading. These examples demonstrate that our generated distracting conditions maintain contextual coherence while effectively increasing reasoning difficulty.
IV-C Performance Drop with the IGC-MWP Dataset
In addition to the qualitative analysis in Sec. IV-B, we quantitatively evaluate LLM performance on the IGC-MWP dataset. Intuitively, well-designed distracting conditions should make problems more challenging, resulting in a greater performance gap relative to the clean dataset.
For comparison, we include two existing GSM-8K variants: GSM-8K-Adv and GSM-Symbolic, both of which augment the original dataset with distracting conditions. As shown in Table II, all GSM variants cause notable performance drops compared to the base GSM-8K. Importantly, the performance degradation on our IGC-MWP dataset is the largest among them, suggesting that our generated problems are of higher quality and present more realistic and cognitively demanding distractions for LLMs.
@@ IV-C Performance Drop with the IGC-MWP Dataset
IGC-GSM 40.24 21.76
TABLE II: Performance and Performance Drop on GSM Variants Datasets. The backbone LLM is Llama2-7B.
V Analysis and Discussion
V-A Case Study
@@
Fail Step       Example (Question and Shared Solution)   Explanation for Failure Reason
Initial
Generation
(Step 1)
@@
Question: Betty is saving money for a new wallet which costs \$100. Betty has only half of the money she needs. Her parents decided to give her \$15 for that purpose, and her grandparents twice as much as her parents. Later, her grandparents added \$2 for her good grades. How much more money does Betty need to buy the wallet? Shared Solution: In the beginning, Betty has only 100 / 2 = \$<<100/2=50>>50. Betty’s grandparents gave her 15 * 2 = \$<<15*2=30>>30. This means, Betty needs 100 - 50 - 30 - 15 = \$<<100-50-30-15=5>>5 more. #### 5 Explanation: NO. The shared solution is inconsistent with the revised question, since the added condition changes how much Betty gets from her grandparents. They now give \$32 (including the extra \$2), so the correct result should be 100-50-15-32=3, not 5. The added condition alters both the reasoning and the final answer.
Quantitative
Relationship
(Step 2)
@@
Question: Lily has 5 lottery tickets to sell. She sells the first ticket for \$1. She then sells each successive ticket for a dollar more than the previous ticket. Each buyer also pays an additional 10% service fee, which Lily donates to charity and does not include in her profit. She plans to keep a \$4 profit and give the remaining money as the prize. How much money will the winner of the lottery receive? Shared Solution: The second ticket sold will cost \$1 + \$1 = \$<<1+1=2>>2. The third ticket sold will cost \$2 + \$1 = \$<<2+1=3>>3. The fourth ticket sold will cost \$3 + \$1 = \$<<3+1=4>>4. The fifth ticket sold will cost \$4 + \$1 = \$<<4+1=5>>5. The total money collected is \$1 + \$2 + \$3 + \$4 + \$5 = \$<<1+2+3+4+5=15>>15. After taking profit, the total prize money will be \$15 - \$4 = \$<<15-4=11>>11. #### 11 Explanation: NO. Although the added condition introduces numerical information (the 3-digit serial number), it does not influence the reasoning process or the arithmetic relationships in the problem. The prices, total revenue, and final prize remain unchanged. Therefore, this new condition is mathematically irrelevant but cognitively distracting—an example of a non-trivial yet irrelevant condition.
Difficulty
Checking
(Step 3)
@@
Question: Tobias is buying a new pair of shoes that costs \$95. He has been saving up his money each month for the past three months. He gets a \$5 allowance a month. He also mows lawns and shovels driveways. He charges \$15 to mow a lawn and \$7 to shovel. Last month, he also helped his neighbor wash a car for \$10, but he decided to save that money for a different purchase. After buying the shoes, he has \$15 in change. If he mows 4 lawns, how many driveways did he shovel? Shared Solution: He saved up \$110 total because 95 + 15 = <<95+15=110>>110 He saved \$15 from his allowance because 3 x 5 = <<3*5=15>>15 He earned \$60 mowing lawns because 4 x 15 = <<4*15=60>>60 He earned \$35 shoveling driveways because 110 - 60 - 15 = <<110-60-15=35>>35 He shoveled 5 driveways because 35 / 7 = <<35/7=5>>5 #### 5 Explanation: NO. Although the new condition introduces a numerical value, it only adds or subtracts simple amounts without affecting reasoning complexity. The distraction is too trivial to mislead solvers or increase cognitive load, making it an ineffective distractor.
Desirable
Traits
(Step 4/5)
@@
Question: Frank needs to meet a quota at work for his sales. It’s the beginning of the month and in 30 days he needs to have 50 cars sold. The first three days he sold 5 cars each day. Then the next 4 days he sold 3 cars each day. The dealership awards 1 bonus point for every 2 cars sold, and Frank currently has 12 bonus points this month. If the month is 30 days long how many cars does he need to sell for the remaining days to meet his quota? Shared Solution: On days one, two, and three he sold 5 cars each day so, 5 cars + 5 cars + 5 cars = <<5+5+5=15>>15 cars that he sold on those days. On days 4,5,6 and 7 he sold 3 cars each day so, 3 cars + 3 cars + 3 cars + 3 cars = <<3+3+3+3=12>>12 cars that he sold on those days. Now we combine the cars he has already sold, which is 15 cars, and 12 cars, so 15 cars + 12 cars = <<15+12=27>>27 cars are sold in total. If he needs to sell 50 cars and has already sold 27 cars then 50 cars - 27 cars = <<50-27=23>>23 cars left to sell to meet his quota. #### 23 Explanation: NO. The added condition is not consistent/reconcilable with Q1: 12 points imply 24 sold cars, conflicting with the stated 27. Although numerically grounded and seemingly relevant, it violates the desirable trait of contextual consistency while leaving the solution unchanged and adding cognitive distraction.
@@
TABLE III: Representative failed cases across different stages of the IGC-MWP framework. Each example includes the shared solution, the revised question (with the red text indicating the added distracting condition—the question without it corresponds to the original version), and the LLM-generated explanation describing the reason for failure.
@@
As shown in Table III, most failures fall into one of the four major categories. The first two types—quantitative relationship and difficulty checking—reflect structural limitations in the generation prompts, while the latter two highlight semantic and contextual challenges in maintaining narrative coherence. Notably, even larger models encounter difficulties in reasoning about implicit contextual logic, underscoring the complexity of multi-step cognitive reasoning required for high-quality MWP generation. These analyses suggest that incorporating explicit reasoning chains and reinforcement from LLM feedback could further enhance the robustness of future frameworks.
VI Conclusion
@@
In this work, we propose an LLM-driven iterative framework, IGC-MWP, for generating distracting conditions in math word problems (MWPs). Unlike traditional human-based or rule-based methods that require extensive manual effort, IGC-MWP leverages large language models to automatically analyze existing problem contexts and generate new distracting conditions with minimal human intervention.

### v3

@@ 3.3 Rejection Mechanism
Regeneration Prompts The question you generate does not comply with all the requirements. Considering the following explanation on your errors, revise the question again. Take the error and initial requirements into consideration. Directly give the full revised question without explanations. Q1: {Original Question} Q2: {Revised Question} Explanation: The solution does not fit the question because the equation used in the solution is incorrect ... should be 5, not 6.
@@
4 Experiment and Result
4.1 Settings
@@ 4.2 Generated Dataset
As illustrated, the problems in the IGC-MWP dataset exhibit clear advantages and avoid the undesirable characteristics discussed earlier. For instance, in Problem 1 of Figure 8, the original story describes slices of pizza, and the added condition naturally follows this theme rather than introducing unrelated context. The original question involves two pizza sizes, while the added condition introduces a medium pizza with a comparable number of slices (12). This new condition adds meaningful distraction—students must now reason about the number of slices in each pizza type and correctly distinguish between large, small, and medium sizes.
Similarly, in Problems 2 and 3 (Figure 8), the added conditions involve the number of lines on pages or the number of pages in books, both of which are conceptually tied to the core entities of the original stories about reading. These examples demonstrate that our generated distracting conditions maintain contextual coherence while effectively increasing reasoning difficulty.
4.3 Performance Drop with the IGC-MWP Dataset
In addition to the qualitative analysis in Sec. 4.2, we quantitatively evaluate LLM performance on the IGC-MWP dataset. Intuitively, well-designed distracting conditions should make problems more challenging, resulting in a greater performance gap relative to the clean dataset.
For comparison, we include two existing GSM-8K variants: GSM-8K-Adv and GSM-Symbolic, both of which augment the original dataset with distracting conditions. As shown in Table 3, all GSM variants cause notable performance drops compared to the base GSM-8K. Importantly, the performance degradation on our IGC-MWP dataset is the largest among them, suggesting that our generated problems are of higher quality and present more realistic and cognitively demanding distractions for LLMs.
@@ 4.3 Performance Drop with the IGC-MWP Dataset
IGC-MWP 40.24 21.76
Table 3: Performance and Performance Drop on GSM Variants Datasets. The backbone LLM is Llama2-7B.
5 Analysis and Discussion
5.1 Case Study
@@
Fail Step Example (Question and Shared Solution) Generated Explanation
@@
      Initial Generation (Step 1) Question: Betty is saving money for a new wallet which costs \$100. Betty has only half of the money she needs. Her parents decided to give her \$15 for that purpose, and her grandparents twice as much as her parents. Later, her grandparents added \$2 for her good grades. How much more money does Betty need to buy the wallet? Shared Solution: In the beginning, Betty has only 100 / 2 = \$<<100/2=50>>50. Betty’s grandparents gave her 15 * 2 = \$<<15*2=30>>30. This means, Betty needs 100 - 50 - 30 - 15 = \$<<100-50-30-15=5>>5 more. #### 5 Explanation: NO. The shared solution is inconsistent with the revised question, since the added condition changes how much Betty gets from her grandparents. They now give \$32 (including the extra \$2), so the correct result should be 100-50-15-32=3, not 5. The added condition alters both the reasoning and the final answer.
@@
      Quantitative Relationship (Step 2) Question: Lily has 5 lottery tickets to sell. She sells the first ticket for \$1. She then sells each successive ticket for a dollar more than the previous ticket. Each buyer also pays an additional 10% service fee, which Lily donates to charity and does not include in her profit. She plans to keep a \$4 profit and give the remaining money as the prize. How much money will the winner of the lottery receive? Shared Solution: The second ticket sold will cost \$1 + \$1 = \$<<1+1=2>>2. The third ticket sold will cost \$2 + \$1 = \$<<2+1=3>>3. The fourth ticket sold will cost \$3 + \$1 = \$<<3+1=4>>4. The fifth ticket sold will cost \$4 + \$1 = \$<<4+1=5>>5. The total money collected is \$1 + \$2 + \$3 + \$4 + \$5 = \$<<1+2+3+4+5=15>>15. After taking profit, the total prize money will be \$15 - \$4 = \$<<15-4=11>>11. #### 11 Explanation: NO. Although the added condition introduces numerical information (the 3-digit serial number), it does not influence the reasoning process or the arithmetic relationships in the problem. The prices, total revenue, and final prize remain unchanged. Therefore, this new condition is mathematically irrelevant but cognitively distracting—an example of a non-trivial yet irrelevant condition.
@@
      Difficulty Checking (Step 3) Question: Tobias is buying a new pair of shoes that costs \$95. He has been saving up his money each month for the past three months. He gets a \$5 allowance a month. He also mows lawns and shovels driveways. He charges \$15 to mow a lawn and \$7 to shovel. Last month, he also helped his neighbor wash a car for \$10, but he decided to save that money for a different purchase. After buying the shoes, he has \$15 in change. If he mows 4 lawns, how many driveways did he shovel? Shared Solution: He saved up \$110 total because 95 + 15 = <<95+15=110>>110 He saved \$15 from his allowance because 3 x 5 = <<3*5=15>>15 He earned \$60 mowing lawns because 4 x 15 = <<4*15=60>>60 He earned \$35 shoveling driveways because 110 - 60 - 15 = <<110-60-15=35>>35 He shoveled 5 driveways because 35 / 7 = <<35/7=5>>5 #### 5 Explanation: NO. Although the new condition introduces a numerical value, it only adds or subtracts simple amounts without affecting reasoning complexity. The distraction is too trivial to mislead solvers or increase cognitive load, making it an ineffective distractor.
@@
      Desirable/Undesirable Traits (Step 4/5) Question: Frank needs to meet a quota at work for his sales. It’s the beginning of the month and in 30 days he needs to have 50 cars sold. The first three days he sold 5 cars each day. Then the next 4 days he sold 3 cars each day. The dealership awards 1 bonus point for every 2 cars sold, and Frank currently has 12 bonus points this month. If the month is 30 days long how many cars does he need to sell for the remaining days to meet his quota? Shared Solution: On days one, two, and three he sold 5 cars each day so, 5 cars + 5 cars + 5 cars = <<5+5+5=15>>15 cars that he sold on those days. On days 4,5,6 and 7 he sold 3 cars each day so, 3 cars + 3 cars + 3 cars + 3 cars = <<3+3+3+3=12>>12 cars that he sold on those days. Now we combine the cars he has already sold, which is 15 cars, and 12 cars, so 15 cars + 12 cars = <<15+12=27>>27 cars are sold in total. If he needs to sell 50 cars and has already sold 27 cars then 50 cars - 27 cars = <<50-27=23>>23 cars left to sell to meet his quota. #### 23 Explanation: NO. The added condition is not consistent/reconcilable with Q1: 12 points imply 24 sold cars, conflicting with the stated 27. Although numerically grounded and seemingly relevant, it violates the desirable trait of contextual consistency while leaving the solution unchanged and adding cognitive distraction.
@@
Table 4: Representative failed cases across different stages of the IGC-MWP framework. Each example includes the shared solution, the revised question, and the LLM-generated explanation describing the reason for failure. The added irrelevant conditions are marked with underlines
@@
As shown in Table 4, most failures fall into one of the four major categories. The first two types—quantitative relationship and difficulty checking—reflect structural limitations in the generation prompts, while the latter two highlight semantic and contextual challenges in maintaining narrative coherence. Notably, even larger models encounter difficulties in reasoning about implicit contextual logic, underscoring the complexity of multi-step cognitive reasoning required for high-quality MWP generation. These analyses suggest that incorporating explicit reasoning chains and reinforcement from LLM feedback could further enhance the robustness of future frameworks.
6 Conclusion
@@
In this work, we propose an LLM-driven iterative framework, IGC-MWP, for generating distracting conditions in math word problems (MWPs). Unlike traditional human-based or rule-based methods [20, 3] that require extensive manual effort, IGC-MWP leverages large language models to automatically analyze existing problem contexts and generate new distracting conditions with minimal human intervention.